OpenStudy (anonymous):

The slope of a line is 16 and the point (6, 18) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

use the general form y-y1 = m(x-x1} where m=slope and x1 and y1 are the point on the line then rearrange to slope-intercept

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\ where\ m=16,\ y_1=18,\ and\ x_1=6\]

OpenStudy (anonymous):

Jeez, I can't paste to save my life....